# AR-1-to-3: Single Image to Consistent 3D Object via Next-View Prediction

### v3

@@ 1 Introduction
Synthesizing 3D objects from a single-view image has been a fundamental and enduring research in both computer vision and graphics communities. This task is challenging, as it not only involves reconstructing the visible regions of a 3D object but also necessitates extrapolating the invisible parts. To address these challenges, researchers have sought to synthesize novel views [31, 4, 25, 22] with the powerful priors of diffusion models, which have achieved remarkable success in image generation [33, 29, 19, 61].
@@
Extensive quantitative and quantitative experiments on large-scale 3D datasets demonstrate that our approach can generate more consistent 2D multi-view images than previous works and produce high-quality 3D assets.
@@ 2 Related Work
Diffusion Models [11, 33, 29, 16] pre-trained on large-scale 2D datasets have demonstrated remarkable performance in generating high-quality images and powerful zero-shot generalization. In recent years, significant effort has been consecutively devoted to transferring such merits of 2D diffusion models to 3D generation.
DreamFusion [30], SJC [49], and Fantasia3D [2] propose to distill the knowledge of a pre-trained 2D diffusion model by feeding the rendered views to it and performing per-shape optimization but suffer from artifacts such as over-saturated colors and the “multi-face” problem. Zero123 [23] pioneers an open-world single-image-to-3D framework, in which diffusion models are fine-tuned to synthesize new views conditioned on an input view and a set of discrete camera poses. ImageDream [50] adopts world camera coordination as in MVDream [37] to recover 3D geometries. Magic123 [31] combines the 3D prior of Zero123 and the 2D prior of the stable diffusion model together to enhance the quality of generated 3D meshes. One-2-3-45 [22] uses Zero123 to generate multi-view images, which are lifted to 3D space to assist in the generation of 3D meshes. Several approaches like Consistent123 [20] and MVD-Fusion [14], improve Zero123 by incorporating additional priors, i.e., boundary and depth. Besides, more and more attention has been drawn to enforce consistency between the generated multiple views. SyncDreamer [24] adopts a 3D-aware attention mechanism to correlate the corresponding features across different views. MVDiffusion [41] generates multi-view images in parallel through the weight-sharing multi-branch UNet with shared weights and correspondence-aware attention. More recently, Cycle3D [42] and Free3D [60] simultaneously generate multiple views to model their joint distribution. Zero123++ [35] further proposes a strategy of tiling six target views surrounding the 3D object into a grid image. This strategy has been successively adopted by follow-up works such as One-2-3-45++ [21], Instant3D [17], and InstantMesh [54].
@@ 2.2 Autoregressive Generation
The autoregressive scheme, commonly employed in time series analysis and prediction, operates on the fundamental principle that the current value in a sequence can be explained by its preceding values. Based upon this technology, researchers have developed a series of classical sequential modeling methods [12, 36, 47, 28]. In recent years, there has been a growing interest in extending such an autoregressive fashion to various communities.
PixelRNN [45] is one of the pioneering methods to generate high-quality images with intricate details by modeling pixel dependencies. VQ-VAE [46] revolutionizes the learning of discrete representations by incorporating the codebook mechanism, enabling efficient encoding and decoding processes of images. VQ-GAN [9] adopts a transformer architecture to model serialized visual parts and introduces adversarial loss during the training process. Parti [55] proposes a pathways autoregressive model treating the image generation as a sequence-to-sequence modeling to generate high-fidelity photorealistic images. VAR [43], LlamaGen [39], and Infinity [10] scale image generation by incorporating multimodal large models. In addition, some approaches [34, 15, 1, 48, 62] attempt to integrate diffusion models with the sequential strategy to achieve more temporally consistent video generation. More recently, MeshGPT [38], MeshAnything [5], and MeshXL [3] proposes to generate triangle faces in an autoregressive manner for artist-created 3D meshes. TAR3D [58] and SAR3D [6] quantify 3D latent representations, which are utilized to generate 3D objects with the next-token prediction strategy.
@@ 5 Conclusions
Supplementary Material
Appendix A More Ablative Results
Ablation on Local Feature Encoding. When the length of the view sequence becomes excessively long, the Stacked-LE strategy may encounter GPU burden issues. Inspired by recent works in video generation [61], we introduce a random sampling design to alleviate this concern. Specifically, we retain all tokens of as they originate from the input image, and perform random sampling with a proportion of on the L tokens from each generated novel view. When the value is set to 0, the Stacked-LE strategy degenerates to the original Reference attention; when is 1, it is equivalent to using all tokens as in Eqn. 2 of the main text.
We investigate the impact of on the performance and efficiency of our model in a train-free fashion. We freeze the trained model parameters and calculate the two measure values, i.e., PSNR and FLOPS, at different sampling rates. As shown in Fig. 9, the PSNR value drops sharply when is 0. We believe that the reason is that the parameters of AR-1-to-3 have been optimized based on additional view features. If the model misses them, ambiguity may arise in understanding 3D objects. When is greater than 0, the PSNR value gradually increases with . Meanwhile, the FLOPS value, which we calculate according to the maximum autoregressive sequence length, continues to increase. To achieve a trade-off between performance and efficiency, it is recommended to take a value between 0.3 and 1 based on the sequence length and GPU resources.
Appendix B More Visualization Results
Apart from the quantitative and qualitative evaluation in the main text, we also provide more visualization examples from the benchmark dataset, i.e., Objaverse [7], and the out-of-domain dataset, i.e., GSO [8] and Omni3D [53], to further validate the effectiveness of our AR-1-to-3.
Novel View Synthesis. We select three types of hard cases , including word-object, humanoid-object, and multi-object, to highlight the superiority of the proposed next-view prediction paradigm for novel view synthesis. As shown in Fig. 10, Fig. 11, and Fig. 12, the novel views progressively generated by our AR-1-to-3 (blue boxes) can capture intricate geometric and textural details of various complex objects, closely mirroring the given images (black boxes).
Image-to-3D Generation. As shown in Fig. 13, our AR-1-to-3 can produce high-quality 3D objects that adhere to diverse and complex image prompts, which should be attributed to the consistent novel views generated by our next-view prediction. These results demonstrate that assigning different generation priorities based on the camera poses difference between novel views and the input view is a promising paradigm towards high-fidelity 3D asset creation.
Appendix C Future Work
This work establishes a next-view prediction paradigm for consistent novel view synthesis and 3D asset creation. However, constrained by computational resources, several promising research directions remain unexplored and are reserved for future work. 1) Scalability. To obtain more powerful 3D generation capabilities, we plan to curate a larger collection of high-quality 3D data from ObjaverseXL and replace our SD diffusion model with DiT. 2) Training Optimization. To better facilitate next-view prediction modeling, we intend to upgrade the supervision from the views of current views to the total view sequence. 2) Method Extension. We will explore to extend the proposed next-view prediction paradigm from Zero123++ to other multi-view diffusion models, e.g., ImageDream [50].

### v4

@@ 1 Introduction
Synthesizing 3D objects from a single-view image has been a fundamental and enduring research in both computer vision and graphics communities. This task is challenging, as it not only involves reconstructing the visible regions of a 3D object but also necessitates extrapolating the invisible parts. To address these challenges, researchers have sought to synthesize novel views [40, 4, 32, 29] with the powerful priors of diffusion models, which have achieved remarkable success in image generation [43, 38, 25, 74, 7].
@@
Extensive quantitative and qualitative experiments on large-scale 3D datasets demonstrate that our approach can generate more consistent 2D multi-view images than previous works and produce high-quality 3D assets.
@@ 2 Related Work
Diffusion Models [43, 38, 20, 8, 11] pre-trained on large-scale 2D datasets have demonstrated remarkable performance in generating high-quality images and powerful zero-shot generalization. In recent years, significant effort has been consecutively devoted to transferring such merits of 2D diffusion models to 3D generation.
Early methods [39, 2, 71, 23, 41] distill knowledge of 2D pre-trained models by feeding the rendered views to it and performing per-shape optimization but suffer from artifacts such as over-saturated colors and the “multi-face” problem. Zero123 [30] pioneers an open-world single-image-to-3D framework, in which diffusion models are fine-tuned to synthesize new views conditioned on an input view and a set of discrete camera poses. ImageDream [60] adopts world camera coordination as in MVDream [47] to recover 3D geometries. Magic123 [40] combines the 3D prior of Zero123 and the 2D prior of the stable diffusion model together to enhance the quality of generated 3D meshes. One-2-3-45 [29] uses Zero123 to generate multi-view images, which are lifted to 3D space to assist in the generation of 3D meshes. Several approaches like Consistent123 [27] and MVD-Fusion [17], improve Zero123 by incorporating additional priors, i.e., boundary and depth. Besides, more and more attention has been drawn to enforce consistency between the generated multiple views. SyncDreamer [31] adopts a 3D-aware attention mechanism to correlate the corresponding features across different views. MVDiffusion [51] generates multi-view images in parallel through the weight-sharing multi-branch UNet with shared weights and correspondence-aware attention. Recent methods [52, 73, 22, 13] explore to simultaneously generate multiple views to model their joint distribution. Zero123++ [45] further proposes a strategy of tiling six target views surrounding the 3D object into a grid image. This strategy has been successively adopted by follow-up works such as One-2-3-45++ [28], Instant3D [21], and InstantMesh [65].
@@ 2.2 Autoregressive Generation
The autoregressive scheme, which operates on the fundamental principle that the current value in a sequence can be explained by its preceding values, plays a vital role in the deep learning era [37, 59, 69, 64, 35]. Based upon this technology, researchers have developed a series of classical sequential modeling methods [15, 46, 57, 36, 26]. In recent years, there has been a growing interest in extending such an autoregressive fashion to various communities.
PixelRNN [55] is one of the pioneering methods to generate high-quality images with intricate details by modeling pixel dependencies. VQ-VAE [56] revolutionizes the learning of discrete representations by incorporating the codebook mechanism, enabling efficient encoding and decoding processes of images. VQ-GAN [12] adopts a transformer architecture to model serialized visual parts and introduces adversarial loss during the training process. Parti [66] proposes a pathways autoregressive model treating the image generation as a sequence-to-sequence modeling to generate high-fidelity photorealistic images. VAR [53], LlamaGen [49], and Infinity [14] scale image generation by incorporating multimodal large models. In addition, some approaches [44, 19, 1, 58, 77] attempt to integrate diffusion models with the sequential strategy to achieve more temporally consistent video generation. More recently, MeshGPT series [48, 5, 3] propose to generate triangle faces in an autoregressive manner for artist-created 3D meshes. TAR3D [70] and SAR3D [6] quantify 3D latent representations [18], which are used to generate SDF-based 3D objects [75, 76] with the next-token prediction strategy.
@@ 5 Conclusions
Acknowledgments. This research was partially funded by Shenzhen Science and Technology Program (No. JCYJ20240813114237048). Besides, we express our gratitude to Dr. Xin Chen for his assistance in revising the paper.